# CRoP: Context-wise Robust Static Human-Sensing Personalization

### v4

@@ 8.1. Pruning mechanism
Magnitude-based ($MP$) vs. Gradient-based Pruning ($GP$): Liu et al. (2020) introduced gradient-based pruning ($GP$), which assigns importance to model parameters (i.e., kernels or nodes) based on the $\ell_{1}$ norm of gradients (computed from the training set data) and prunes the least important ones. In literature (e.g., (Krizhevsky and Hinton, 2009)), the approach has been shown to work on models trained and tested on the data having independent and identical distribution (IID). However, in this work, data from contexts $\mathcal{C}_{a}$ and $\mathcal{C}_{u}$ follow different distributions. Since, according to the problem setup, only $\mathcal{C}_{a}$ data is accessible, using GP in CRoP leads to overfitting on $\mathcal{C}_{a}$ and poor performance on unseen contexts $\mathcal{C}_{u}$. In contrast, $MP$ selects important parameters based solely on weight magnitude, making it more robust and resulting in better performance on $\mathcal{C}_{u}$. As shown in Figures 7(a), $GP$ performs well on $\mathcal{C}_{a}$ but consistently underperforms compared to $MP$ on $\mathcal{C}_{u}$ across all three users.
Figure 7. Ablation study exploring different alternatives for (a) pruning and (b) regularization mechanisms using inference accuracy as the metric of evaluation
Top prune ($MP-T$) vs. lower prune ($MP$): CRoP’s conventional magnitude-based pruning ($MP$) removes a percentage of the lowest-magnitude weights, which typically have minimal impact on model inference (Luo et al., 2017; Zhu and Gupta, 2017). However, Bartoldson et al. (2020) suggests that pruning the top weights and retraining them ($MP-T$) can improve generalization by creating flatter loss landscapes. Although $MP-T$ yields models that generalize better than traditional personalized models, Figure 7(a) shows that it performs significantly inferior to CRoP for all users in both contexts. This is because CRoP aims to isolate a user-specific sub-network that performs well in the available context $\mathcal{C}_{a}$, rather than finding a generic sub-network. Pruning top weights removes user-specific information learned in step 2 of the Algorithm 1, which slightly improves performance in unseen context $\mathcal{C}_{u}$ but significantly degrades performance in context $\mathcal{C}_{a}$. Moreover, since the removed top weights are important for $\mathcal{C}_{a}$, the final finetuning stage (step 5 of the Algorithm 1) recaptures that information, overwriting the generic information, nullifying the approach’s effect.
@@
In the one-shot approach, the ‘ToleratedPrune’ module (Algorithm 1 line 3 and Algorithm 2) calculates the optimal pruning amount by removing the maximum parameters without exceeding the accuracy loss threshold $\tau$ for context $\mathcal{C}_{a}$. In the iterative approach, we gradually reach the same pruning amount by removing small fractions of parameters for each cycle. These parameters are replaced by weights from the generic model, but since the finetuning loss function resembles the initial one, it drives those weights to lower magnitudes, causing them to be pruned again in the next cycle. This repeats until the target pruning amount is reached, effectively pruning a similar set of weights as the one-shot method but in smaller steps.
@@ 9. Run-time analysis
CRoP is a static personalization approach that requires one-time on-device training of the generic model during personalization. We performed run-time evaluations to assess the viability of the CRoP framework across five different deployment platforms: AMD Ryzen 9 5950X 16-Core Processor, AMD Ryzen 9 7950X 16-Core Processor, Apple M1, NVIDIA RTX A6000 and NVIDIA GeForce RTX 4090. Table 6 shows the resources required in the training phase in terms of training time (s), process memory (MBs), and resource, i.e., GPU/CPU utilization (%) for two health-related datasets: PERCEPT-R and Stress-Sensing. According to our evaluation, except for the scalable device, Apple M1, the training time is a couple of seconds, and resource requirements are minimal. Even on scalable devices like Apple M1, computation time remains efficient, with only a slight increase, typically adding at most a few minutes. Since this is a one-time process, CRoP is efficient compared to continuous learning methods that require repeated adaptation. Particularly, this makes CRoP practical for critical applications, such as clinical settings, where it incurs only a brief, one-time computation overhead during enrollment.
Dataset Platform Train (s) RSS (MBs) Train Usage (%)
PERCEPT-R AMD Ryzen 9 5950X 16-Core Processor 15.18 2787.36 24.99
Apple M1 102.39 352 0.58
NVIDIA RTX A6000 1.94 1132.61 20
AMD Ryzen 9 7950X 16-Core Processor 14.22 830.97 24.9
NVIDIA GeForce RTX 4090 1.54 1279.2 36
Stress-Sensing AMD Ryzen 9 5950X 16-Core Processor 6.95 1469.59 24.9
Apple M1 11.71 77 0.85
NVIDIA RTX A6000 8.49 1071.14 7
AMD Ryzen 9 7950X 16-Core Processor 3.13 633.27 25
NVIDIA GeForce RTX 4090 4.31 1249.79 14
Table 6. Resource requirement for one time training using CRoP
10. Limitations and Future Direction
Some limitations and future research are discussed below:
  1. (1)
@@ 10. Limitations and Future Direction
    The paper performed a limited evaluation on pruning paradigms through ablation studies as it was not the primary focus of the study. Section 8 on ablation study justifies CRoP’s design choice but does not establish any particular paradigm’s superiority in unseen contexts.
  2. (2)
@@
    As the approach relies on using a pre-trained off-the-shelf model as an input, the quality of this model can impact the performance of the final personalized models. As discussed in Section 6.3.3, users with suboptimal performance of the generic model show higher gain through personalization. On the other hand, for certain participants in the PERCEPT-R dataset, a high-performing model leaves minimal room for improvement through personalization.
  3. (3)
@@
    We restrict our study to the models benchmarked and deployed for datasets used in this work without accounting for model variability. However, the different model architectures employed across all the datasets incorporate a range of layers such as convolutional, linear, BiLSTM, and GRU, which introduces some variability.
  4. (4)
    Some evaluation results, for instance, Table 4 evaluations are shown as average. However, the metrics $\Delta_{P}$ and $\Delta_{G}$ are calculated individually for each person. This is because models personalized for one user cannot be applied to other users in real-world situations. Consequently, we focus our evaluations on intra-user generalizability, excluding discussion for inter-user or inter-dataset generalizability.
11. Broader Impact
This paper addresses a critical research gap, enhancing the practical utility of human-sensing solutions in real-world applications, particularly in automated healthcare. Next-generation healthcare systems, which employ neural networks for tasks ranging from daily activity detection (Ustev et al., 2013; Stisen et al., 2015) to safety-critical conditions like atrial fibrillation (Comstock, 2017), benefit from personalization due to the heterogeneity in health sensing data (Ji et al., 2021; Sempionatto et al., 2021). CRoP offers several advantages:
  1. (1)
    Eliminating the Generic Model Training: CRoP leverages off-the-shelf pre-trained models, eliminating the need for training generic models. It is especially valuable in clinical settings where privacy concerns restrict data sharing for training purposes (Malin et al., 2018). This increases the feasibility of deploying personalized models in healthcare.
  2. (2)
    No privacy concerns: CRoP operates on local devices, eliminating the need for transferring potentially sensitive information to a central server. To demonstrate scalability on local devices, the resource consumption for CRoP personalization on 5 devices is shown in Section 9.
  3. (3)
    Flexibility to use any model architectures: As model pruning has proven applicable to various model architectures, CRoP is not restricted to any model architecture constraints, ensuring wide-ranging applicability.
@@ 12. Conclusion
This study introduces CRoP, a novel static personalization approach generating context-wise intra-user robust models from limited context data. Using pruning in a novel way to balance personalization and generalization, empirical analysis on four human-sensing datasets shows CRoP models exhibit an average increase of $35.23\%$ in personalization compared to generic models and $7.78\%$ in generalization compared to conventionally-finetuned personalized models. CRoP utilizes off-the-shelf models, reducing training effort and addressing privacy concerns. With practical benefits and quantitative performance enhancements, CRoP facilitates reliable real-world deployment for AI-based human-sensing applications like healthcare.
@@ Appendix A Ensuring Reproducibility
A.1. Metrics for classification accuracy evaluation
@@
A.2. Hyperparameters
The approach uses several hyperparameters for generic model training and personalization. Table 7 and 8 show the hyperparameter values for generic and personalized model training, respectively. These values correspond to the best results obtained for the data belonging to the available context using a grid search. For training the generic model, in addition to the number of epochs, ‘Base Learning Rate’ and ‘Max Learning Rate’ (the arguments for CycleLR [Smith, 2017]) are the hyperparameters. For the personalized model, learning rate (fixed), $\alpha$, $\tau$, number of epochs for initial finetuning (Initial Epochs), and epochs for final finetuning (Final Epochs) are the hyperparameters. The range of these hyperparameters used for grid search during personalization is also mentioned in Table 8.
@@ A.2. Hyperparameters
A.3. Code
The code is provided at anonymous link arranged into dataset-specific folders. Each folder contains the pre-trained generic model, all the required modules, and the instructions to run the code. The seed values used for the evaluations are also provided in the shell files. The data partitioned into personalized and context-wise sets will be released upon publication.
A.4. Compute Resources
@@ B.1. Detailed discussion of $\Delta_{P}$ results
The physiological features used in this dataset vary significantly from one user to other. Thus, Tables 8(e)-8(h) show that the generic models do not perform well on personalized data. Personalized finetuning enables the model to learn person-specific patterns, allowing the model’s performance to improve not only in the available context but also in the unseen context. This results in average personalization benefit $(\Delta_{P})$ of $67.81$ and $85.25$ for Scenario 1 and Scenario 2, respectively. It is important to note that for each Scenario, only one model is trained for the available context and tested for two different unseen contexts. Moreover, double context change (Tables 8(g) and 8(h)) shows lower personalization benefit as compared to single context change (Tables 8(e) and 8(f)).
@@ B.2. Detailed discussion of $\Delta_{G}$ results
As observed in Tables 8(e)-8(h), personalized finetuning improves models performance on unseen context as well, we can claim that there is some person-specific traits which are common in available and unseen context. While comparing our final models with conventionally-finetuned models (Tables 9(e)-9(h)), performance boost in both available and unseen context could be observed. This can be attributed to the generalization improvement benefits of model pruning [Jin et al., 2022]. This results in average generalization benefit $(\Delta_{G})$ of $13.81$ and $13.08$ for Scenario 1 and Scenario 2, respectively, for single context change. Similar personalization benefits could be seen for double context change.

### v5

@@ 8.1. Pruning mechanism
Figure 7. Ablation study exploring different alternatives for (a) pruning and (b) regularization mechanisms using accuracy as the metric of evaluation
Magnitude-based ($MP$) vs. Gradient-based Pruning ($GP$): Liu et al. (2020) introduced gradient-based pruning ($GP$), which prunes parameters based on the $\ell_{1}$-norm of gradients. While effective for IID data ((Krizhevsky and Hinton, 2009)), in CRoP, where $\mathcal{C}_{a}$ and $\mathcal{C}_{u}$ have different distributions, $GP$ overfits to $\mathcal{C}_{a}$ and performs poorly on $\mathcal{C}_{u}$. In contrast, $MP$ selects parameters by weight magnitude, making it more robust and better performing on $\mathcal{C}_{u}$. As shown in Figures 7(a), $GP$ performs well on $\mathcal{C}_{a}$ but consistently underperforms compared to $MP$ on $\mathcal{C}_{u}$ across all three users.
Top prune ($MP-T$) vs. lower prune ($MP$): Bartoldson et al. (2020) suggests that pruning the top weights and retraining them ($MP-T$) can improve generalization by flattening loss landscapes. However, Figure 7(a) shows that it performs significantly inferior to CRoP for all users in both contexts. This is because CRoP aims to isolate a user-specific sub-network that performs well in the available context $\mathcal{C}_{a}$, while $MP-T$ removes important user-specific information. Moreover, the final finetuning stage (Algorithm 1 Line 5) retrains those pruned parameters highly important for $\mathcal{C}_{a}$, overwriting the generic information impacting context $\mathcal{C}_{u}$.
@@
In the one-shot approach, the ‘ToleratedPrune’ module (Algorithm 1 line 3 and Algorithm 2) calculates the optimal pruning amount by removing the maximum parameters without exceeding the accuracy loss threshold $\tau$ for context $\mathcal{C}_{a}$. In the iterative approach, we gradually reach the same pruning amount by removing small fractions of parameters for each cycle. These parameters are replaced by weights from the generic model, but since the finetuning loss function resembles the initial one, where no early stopping mechanism is used, it drives those weights to lower magnitudes, causing them to be pruned again in the next cycle. This repeats until the target pruning amount is reached, effectively pruning a similar set of weights as the one-shot method but in smaller steps.
@@ 8.2. Regularization Mechanisms
Dataset Platform Train (s) RSS (MBs) Resource Usage (%)
PERCEPT-R AMD Ryzen 9 5950X 16-Core Processor 17.68 2881.13 23
Apple M1 119.56 288 1
NVIDIA RTX A6000 5.16 1108.65 16
AMD Ryzen 9 7950X 16-Core Processor 16.82 807.58 24
NVIDIA GeForce RTX 4090 2.34 1166.82 23
Google Pixel 6 824.19 360.95 36.8
Cortex-A76 (Raspberry Pi 5) 190.44 505.67 49
Stress-Sensing AMD Ryzen 9 5950X 16-Core Processor 8.77 1510.28 21
Apple M1 12.93 134 1
NVIDIA RTX A6000 11.34 1081.73 3
AMD Ryzen 9 7950X 16-Core Processor 3.68 624.98 23
NVIDIA GeForce RTX 4090 5.01 1136.69 16
Google Pixel 6 98.45 246.06 33
Cortex-A76 (Raspberry Pi 5) 50.86 392.74 49
Table 7. Resource requirement for one time training using CRoP
Model $\rightarrow$ Generic Model Personalized Model
Dataset Platform Inference (s) RSS (MBs) Resource Usage (%) Inference (s) RSS (MBs) Resource Usage (%)
PERCEPT-R AMD Ryzen 9 5950X 16-Core Processor 0.03 2752 16 0.03 2752 16
Apple M1 0.22 362.34 0.19 0.22 361 0.19
NVIDIA RTX A6000 0.09 773.33 3 0.09 778.25 1
AMD Ryzen 9 7950X 16-Core Processor 0.04 697 12 0.02 697 12
NVIDIA GeForce RTX 4090 0.07 829.5 0.01 0.07 829.75 0.01
Google Pixel 6 0.62 249.49 37.85 0.61 253.89 38.08
Cortex-A76 (Raspberry Pi 5) 0.2 392.01 49.84 0.18 392.57 49.8
Stress-Sensing AMD Ryzen 9 5950X 16-Core Processor 0.03 510 23 0.03 510 23
Apple M1 0.1 206 0.4 0.1 265 0.4
NVIDIA RTX A6000 0.1 752 2 0.1 753 1
AMD Ryzen 9 7950X 16-Core Processor 0.01 537.14 22.7 0.01 537 22.7
NVIDIA GeForce RTX 4090 0.01 800 1 0.01 800 1
Google Pixel 6 0.13 201.24 25.62 0.11 201.4 29.03
Cortex-A76 (Raspberry Pi 5) 0.14 335.77 48.78 0.15 335.83 48.7
Table 8. Runtime Resource requirement comparison for generic and personalized models obtained for CRoP
@@ 9. Run-time analysis
CRoP is a static personalization approach that requires one-time on-device training of the generic model during personalization. We performed run-time evaluations to assess the viability of the CRoP framework across seven different deployment platforms: AMD Ryzen 9 5950X 16-Core Processor, AMD Ryzen 9 7950X 16-Core Processor, Apple M1, NVIDIA RTX A6000, NVIDIA GeForce RTX 4090, and resource-constrained devices Google Pixel 6 and Cortex-A76 (Raspberry Pi 5). Table 8 shows the average resources required in the training phase in terms of training time (s), process memory (MBs), and resource, i.e., GPU/CPU utilization (%) for two health-related datasets: PERCEPT-R and Stress-Sensing. Notably, the resource utilization reported in Tables 8 and 8 refers to GPU utilization for the NVIDIA RTX A6000 and NVIDIA GeForce RTX 4090 platforms, whereas for all other devices, it reflects CPU utilization. According to our evaluation, except for the resource-constrained device, Apple M1, Google Pixel 6 and Raspberry Pi 5, the training time is a couple of seconds, and resource requirements are minimal. Even on resource-constrained devices like Apple M1, Google Pixel 6 and Raspberry Pi 5, computation time remains efficient, with only a slight increase, typically adding at most a few minutes. Since this is a one-time process, CRoP is efficient compared to continuous learning methods that require repeated adaptation. Particularly, this makes CRoP practical for critical applications, such as clinical settings, where it incurs only a brief, one-time computation overhead during enrollment.
Further, Table 8 shows runtime comparison of the generic off-the-shelf model and the personalized model achieved using CRoP accross seven devices mentioned above . It can be observed that inference time and resource usage for both model states are similar on all devices as the model architecture of the personalized model remains same as that of the off-the-shelf generic model. Thus, CRoP does not introduce any inference time overheads.
10. Discussion, Limitations and Future Direction
Analyzing the pruning pattern across all layers of the model. We employ regularization (during the initial finetuning step) and pruning (during ToleratedPruned step) to identify model parameters that are essential to represent user-specific traits present in the available context $\mathcal{C}_{a}$ data. The remaining parameter values are restored from the generic model to incorporate generic knowledge.
Our analysis of the LeNet model trained on the WIDAR dataset revealed that the initial fine-tuning step has minimal impact on the initial feature extraction layers, as they primarily capture generic low-level features. In contrast, the deeper layers—particularly the last convolutional and first linear layers—undergo significant changes as they learn high-level feature representations essential for user-specific traits. Additionally, the regularization imposed by initial finetuning constrains these learned parameters to a small subset of the deeper layers, enabling the ToleratedPrune module to prune these layers more aggressively than the initial ones. Consequently, CRoP preserves low-level generic features by minimally affecting early layers during initial finetuning while allowing deeper layers to adapt to individual traits. Simultaneously, it integrates more generic information from the generic model during model mixing on those deeper layers by performing more aggressive pruning, ensuring a balance between personalization and generalization. The detailed discussion in provided in Appendix C.
@@ 10. Discussion, Limitations and Future Direction
Evaluating different pruning paradigms The paper performed a limited evaluation on pruning paradigms through ablation studies as it was not the primary focus of the study. Section 8 on ablation study justifies CRoP’s design choice but does not establish any particular paradigm’s superiority in unseen contexts.
@@
Dependence on quality of off-the-shelf generic model. As the approach relies on using a pre-trained off-the-shelf model as an input, the quality of this model can impact the performance of the final personalized models. As discussed in Section 6.3.3, users with suboptimal performance of the generic model show higher gain through personalization. On the other hand, for certain participants in the PERCEPT-R dataset, a high-performing model leaves minimal room for improvement through personalization.
@@
Exploring different model architectures for each dataset. We restrict our study to the models benchmarked and deployed for datasets used in this work without accounting for model variability. However, the different model architectures employed across all the datasets incorporate a range of layers such as convolutional, linear, BiLSTM, and GRU, which introduces some variability.
Assessing the need for inter-user generalizability. The inter-user heterogeneity (Gong et al., 2023; Meegahapola et al., 2023; Sempionatto et al., 2021) kindles the need for personalization. To address this, several human sensing applications (Wang et al., 2016; Burns et al., 2022; Duan et al., 2023; Liu et al., 2022) adopt personalization to generate models tailored for user-specific traits. Notably, Wang et al. (2016) have shown that models trained for a set of users show higher mean average error for data belonging to new users due to inter-user heterogeneity. This challenge parallels domain adaptation scenarios, where each individual can be viewed as a distinct domain, and models adapted to one domain do not transfer well to other domains (Tzeng et al., 2017). Addressing the adaptation of a personalized model to a new user—i.e., inter-user generalizability—remains an open research question. In this work, we restrict our evaluations to the conventional definition of personalization and leave inter-user generalization for future exploration. Additionally, exploring the incremental addition of new users and comparing them against existing incremental learning methods is an interesting direction for future work. This would also allow evaluation of method’s ability to mitigate catastrophic forgetting, which is crucial for broader applicability.
Trade-off between personalized model’s performance in available $\mathcal{C}_{a}$ and unseen $\mathcal{C}_{u}$ context. The goal of personalization in our work extends beyond improving performance in just the available context $\mathcal{C}_{a}$; it aims to enhance the overall performance of the model across all contexts ( that is, both available $\mathcal{C}_{a}$ and unseen $\mathcal{C}_{u}$ contexts). The conventional personalization approaches used as baselines in this work, often update the model using available context data $\mathcal{C}_{a}$ to the extent that results in significant performance drops in unseen contexts $\mathcal{C}_{u}$ and thus, limit the overall performance gains. In contrast, CRoP achieves better balance by improving performance in the available context $\mathcal{C}_{a}$ while minimizing performance degradation in unseen contexts $\mathcal{C}_{u}$. This advantage is formalized in our problem statement and evaluation metric Personalization ($\Delta_{P}$) and Generalization ($\Delta_{G}$). Table 3 reports consistently positive $\Delta_{P}$ and $\Delta_{G}$ values achieved by CRoP for all datasets under all scenarios, indicating that CRoP-personalized models $\mathcal{M}^{P^{a}_{i}}_{\theta}$ achieve overall performance improvements over the generic model $\mathcal{M}^{G}_{\theta}$ and conventionally-finetuned model $\mathcal{M}^{C^{a}_{i}}_{\theta}$, respectively.
@@ 11. Conclusion
This study introduces CRoP, a novel static personalization approach generating context-wise intra-user robust models from limited context data. Using pruning in a novel way to balance personalization and generalization, empirical analysis on four human-sensing datasets shows CRoP models exhibit an average increase of $35.23\%$ in personalization compared to generic models and $7.78\%$ in generalization compared to conventionally-finetuned personalized models. CRoP utilizes off-the-shelf models, reducing training effort and addressing privacy concerns. Since no steps rely on specific model components, CRoP is architecture-agnostic, ensuring broad applicability. With practical benefits and quantitative performance enhancements, CRoP facilitates reliable real-world deployment for AI-based human-sensing applications like healthcare.
Acknowledgements
This work was partly supported by NSF IIS SCH #2124285 and NIH R01 #R01DC020959.
@@ Appendix A Ensuring Reproducibility
A.1. Model Architectures
For each dataset, we employ models used in the literature:
PERCEPT-R
: In line with the literature Benway et al. [2023], we tried several model architectures such as CNN, DNN, BILSTM, etc, whose number of parameters were identified using grid search. Among those, the biLSTM model containing 4 bidirectional LSTM layers followed by 5 linear layers, accompanied by a Hardswish activation layer, was identified as the one that exhibited the best results for the generic data and was used for this study.
WIDAR:
The model used for WIDAR follows the LeNet architecture [Zhang et al., 2022], which contains three 2D convolutional layers followed by two linear layers. Each of these layers, except the final classification layer, is followed by a ReLU activation layer.
ExtraSensory:
The model follows a CNN-GRU-based architecture used in HAR literature [Gong et al., 2019, Hao et al., 2022]. The model consists of three batch-normalized 1D convolution layers followed by a linear layer that feeds into a batch-normalized recursive (GRU) layer and two linear layers to generate embeddings. For the classification head, two linear layers were used.
Stress Sensing Dataset:
The model uses a simple multi-layer-perceptron (MLP) architecture [Eren and Navruz, 2022] consisting of 3 linear layers with hidden size of 128.
A.2. Metrics for classification accuracy evaluation
@@
A.3. Hyperparameters
The approach uses several hyperparameters for generic model training and personalization. Table 9 and 10 show the hyperparameter values for generic and personalized model training, respectively. These values correspond to the best results obtained for the data belonging to the available context using a grid search. For training the generic model, in addition to the number of epochs, ‘Base Learning Rate’ and ‘Max Learning Rate’ (the arguments for CycleLR [Smith, 2017]) are the hyperparameters. For the personalized model, learning rate (fixed), $\alpha$, $\tau$, number of epochs for initial finetuning (Initial Epochs), and epochs for final finetuning (Final Epochs) are the hyperparameters. The range of these hyperparameters used for grid search during personalization is also mentioned in Table 10.
@@ A.3. Hyperparameters
A.4. Code
The code is provided at this Git Repository arranged into dataset-specific folders. Each folder contains the pre-trained generic model, all the required modules, and the instructions to run the code. The seed values used for the evaluations are also provided in the shell files. The data partitioned into personalized and context-wise sets will be released upon publication.
A.5. Compute Resources
@@ B.1. Detailed discussion of $\Delta_{P}$ results
The physiological features used in this dataset vary significantly from one user to other. Thus, Tables 8(e)-8(h) show that the generic models do not perform well on personalized data. Personalized finetuning enables the model to learn user-specific patterns, allowing the model’s performance to improve not only in the available context but also in the unseen context. This results in average personalization benefit $(\Delta_{P})$ of $67.81$ and $85.25$ for Scenario 1 and Scenario 2, respectively. It is important to note that for each Scenario, only one model is trained for the available context and tested for two different unseen contexts. Moreover, double context change (Tables 8(g) and 8(h)) shows lower personalization benefit as compared to single context change (Tables 8(e) and 8(f)).
@@ B.2. Detailed discussion of $\Delta_{G}$ results
As observed in Tables 8(e)-8(h), personalized finetuning improves models performance on unseen context as well, we can claim that there is some user-specific traits which are common in available and unseen context. While comparing our final models with conventionally-finetuned models (Tables 9(e)-9(h)), performance boost in both available and unseen context could be observed. This can be attributed to the generalization improvement benefits of model pruning [Jin et al., 2022]. This results in average generalization benefit $(\Delta_{G})$ of $13.81$ and $13.08$ for Scenario 1 and Scenario 2, respectively, for single context change. Similar personalization benefits could be seen for double context change.